OpenStudy (anonymous):

What is the y coordinate of the parabola with the equation y=x^2-6x+14?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

x coordinate = -b/2a = -(-6)/2(1) = 6/2 = 3 y coordinate = (3)^2 - 6(3) + 14 = 9 -18 + 14 = -9 + 14 = 5